# FactSelfCheck: Fact-Level Black-Box Hallucination Detection for LLMs

### v2

@@ 4.5 Role of Fact-Level Detection in Hallucination Correction
Subsequently, we evaluated the factuality of the corrected responses using the LLM-as-judge approach Zheng et al. (2023). For each corrected sentence, we provided the external biography from Wikipedia as a reference. The reliability of this approach is high, and commonly used in the literature Thakur et al. (2025). We instructed LLM-as-judge to return ’yes’ if the source supported the sentence, ’no’ if it was not, or ’refused’ if the LLM declined to correct the sentence (e.g., due to insufficient knowledge). We then categorized the responses into three labels: ’factual’, ’non-factual’, ’refused’. As mentioned in Section 4.3, we utilized GPT-4o for correction and judging instead of Llama-3.1-70B-Instruct (used in detection). This choice was motivated by the challenging nature of the correction task – the model needs to correct hallucinations using only its internal knowledge, without access to external references. While the model knows hallucinated parts, it must rely on its knowledge to determine the correct information.
@@ 5 Results
This section presents the results of the experiments described in Section 4. The results of the fact-level evaluation are presented in Appendix C.
@@ 5.1 Sentence-Level Detection
$\mathcal{H}$) - 82.56
[…]
$\mathcal{H}$) - 79.02
[…]
@@
Tables 1 and 2 present a comparative analysis of our method against baselines. For WikiBio, FactSelfCheck-Text utilizing $max$ as an aggregation function achieves an AUC-PR score of $92.45$. It demonstrates that our approach is comparable in performance to the leading SelfCheckGPT (SCGPT) variants – Prompt ($93.60$) and NLI ($92.50$). Notably, while our method operates at a more granular level, it maintains competitive performance with a marginal decrease of $1.2\%$ compared to the best SCGPT. It is important to note that comparing our method to SelfCheckGPT at the sentence level inherently disadvantages our approach, as we operate at a more granular level of analysis. Nevertheless, our method still achieves competitive performance despite this inherent challenge.
@@
For FavaMultiSamples, FactSelfCheck-KG (Frequency-based) with max aggregation achieves the highest AUC-PR score of $48.52$, outperforming all sentence-level baselines, including SelfCheckGPT (Prompt) at $46.91$. This result, together with the findings from WikiBio, highlights that the best-performing FactSelfCheck (FSC) variant depends on the dataset. On WikiBio, FSC-Text, which performs direct comparison between facts and samples, consistently achieves the highest hallucination AUC-PR ($92.45$), outperforming FSC-KG, which relies on knowledge graph comparisons. FSC-Text offers computational advantages by eliminating the knowledge graph extraction step. However, on FavaMultiSamples, the frequency-based FSC-KG variant surpasses both FSC-Text and LLM-based FSC-KG. These differences suggest that the optimal variant is influenced by dataset characteristics, such as fact density, sentence length, and text style (e.g., the prevalence of lists in FavaMultiSamples). FSC-KG may be more computationally efficient for longer samples with lower fact density due to reduced token usage, while FSC-Text is preferable for shorter or denser samples. Regarding aggregation functions, $max$ consistently yields better results than $mean$, as it is more sensitive to the presence of hallucinated facts in the text.
@@
An interesting side observation is that our reproduced SCGPT (Prompt) with Llama-3.1-70B-Instruct marginally surpassed the original implementation using GPT-3.5-turbo ($93.60$ vs $93.42$  [^7] The scores on GPT-3.5-turbo were obtained from the original paper Manakul et al. (2023).).
@@ 6 Conclusion
We introduced FactSelfCheck, a novel approach to hallucination detection that operates at the fact-level rather than the sentence or passage level. Our evaluation demonstrates that this fine-grained approach achieves competitive performance with SelfCheckGPT, a less granular method. On the sentence-level detection task, FactSelfCheck achieves AUC-PR scores of $92.45$ and $48.52$ on WikiBio and FavaMultiSamples, respectively, compared to SelfCheckGPT’s $93.60$ and $46.91$. While maintaining comparable performance, our method provides more detailed insights into the nature of hallucinations.
Our fact-level approach improves hallucination correction. Providing the correction system with a list of incorrect facts rather than incorrect sentences allows for more effective corrections. Compared to baseline, we achieved $35.5\%$ increase in factual content and $12.5\%$ reduction in non-factual content. In contrast, the sentence-level approach achieved a $10.6\%$ increase and $4.8\%$ reduction, respectively. It underscores the importance of fine-grained hallucination detection in enhancing the reliability and trustworthiness of systems that use LLMs, particularly in applications requiring high factual accuracy. Finally, we contributed FavaMultiSamples, a novel dataset for evaluating sampling-based methods.
@@
Our study faces two primary limitations. First, we are constrained by the availability of suitable datasets. Although we contributed FavaMultiSamples, the second dataset for evaluating sampling-based methods, it is still a dataset with annotations at the sentence-level. The lack of datasets with long generated passages and annotations at the fact-level forced us to evaluate our method through aggregation rather than directly assessing our fact-level detection capabilities. Second, while the granular approach of FactSelfCheck justifies its increased complexity, the multiple LLM-based steps make it more computationally intensive compared to more straightforward methods like SelfCheckGPT.
@@ Acknowledgments
This work was funded by the European Union under the Horizon Europe grant OMINO – Overcoming Multilevel INformation Overload (grant number 101086321, http://ominoproject.eu/). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. It was also co-financed with funds from the Polish Ministry of Education and Science under the programme entitled International Co-Financed Projects, grant no. 573977. This work was co-funded by the National Science Centre, Poland under CHIST-ERA Open & Re-usable Research Data & Software (grant number 2022/04/Y/ST6/00183).
Appendix A Fava-Sampling Dataset
We built FavaMultiSamples upon the FAVA dataset [^8] https://huggingface.co/datasets/fava-uw/fava-data license: CC-BY-4.0 developed by Mishra et al., 2024. The original FAVA dataset contains $460$ passages generated by GPT (gpt-3.5-turbo-0301) and Llama2-Chat-70B in response to diverse information-seeking prompts. Each passage was annotated by trained annotators for factual accuracy. For more details about the dataset construction and annotation process, please refer to the original paper.
@@ Appendix B Dataset Statistics
Appendix C Fact-Level Detection Evaluation
While the evaluation in Section 5.1 focused on the sentence-level detection due to dataset constraints, we present a complementary evaluation at the fact level here. Since this evaluation is not based on human annotations, we have placed it in Appendix.
Evaluation Data
To enable fact-level evaluation, we annotated facts from the extracted response knowledge graph ($KG_{p}$) using the LLM-as-judge approach Zheng et al. (2023). Since we pass the external Wikipedia biography to the LLM, the reliability of this approach is high, and commonly used in the literature Thakur et al. (2025). This approach annotated whether each fact is supported by the external Wikipedia biography. As a result, we obtained $5488$ binary annotated facts.
Experimental Setup
Results
Method AUC-PR
FactSelfCheck-Text 93.41
FactSelfCheck-KG (LLM-based) 92.25
FactSelfCheck-KG (Freq.-based) 87.99
SelfCheckGPT (Prompt) 86.18
RandomFact 65.79
Table 5 presents the comparative results. FactSelfCheck-Text demonstrates superior performance with an AUC-PR score of $93.41$, followed by FactSelfCheck-KG (LLM-based) at $92.25$. The frequency-based method achieves $87.99$, while SelfCheckGPT (Prompt) scores $86.18$. These results demonstrate the effectiveness of our method in detecting hallucinations at the fact level. Furthermore, the lower performance of averaging the sentence-level scores highlights the importance of designing fact-level methods and validating our approach.
Comparing FactSelfCheck to SelfCheckGPT could be seen as unfair because SCGPT operates at a lower granularity than required by the evaluation task. However, this situation is analogous to the sentence-level evaluation presented in Section 5.1. In both cases, direct comparisons are not fully appropriate. The key difference is that, at the fact level, FactSelfCheck significantly outperforms SelfCheckGPT, while at the sentence level, FactSelfCheck remains competitive despite the unfavorable comparison.
Appendix D Additional Results
D.1 Sentence-Level Detection
@@
$\mathcal{H}$) - 82.56 41.80 62.18
[…]
$\mathcal{H}$) - 79.02 49.18 64.10
[…]
@@
D.2 Effect of Sample Size on Detection Performance
@@
Figure 5 shows a different pattern for FavaMultiSamples, where hallucination detection performance decreases with more samples. This occurs because methods with few samples tend to overestimate hallucination scores, interpreting normal variations as potential hallucinations. As sample size increases, methods become more conservative in their scoring, leading to better calibration. This pattern likely stems from FavaMultiSamples having shorter sentences with fewer facts compared to WikiBio (see Section D.3).
@@
D.3 Evaluation of Intermediate Steps
Entities/ Relations/ Facts/
@@ D.3 Evaluation of Intermediate Steps
Appendix E Computional complexity
@@
Appendix F Case Study of FactSelfCheck vs SelfCheckGPT
@@ Appendix G SelfCheckGPT with Enhanced Prompt
Table 10: Comparison of fact-level FactSelfCheck with sentence-level SelfCheckGPT. An external Wikipedia biography is provided to analyse the correctness of the methods. The red value indicates hallucinations, and the green value indicates factual correctness. The facts were classified using a threshold of $0.4$ utilizing FactSelfCheck, and the sentences were classified using a threshold of $0.75$ with SelfCheckGPT. These thresholds achieved the highest F1-scores in fact-level (Appendix C) and sentence-level (Section 5.1) evaluation, respectively.
Figure 6: Example of a knowledge graph extracted from a response. Edge width represents the hallucination score for each fact, with red edges indicating hallucinated facts and green edges indicating correct facts. Facts were classified using a threshold of $0.4$, which achieved the highest F1-score in the fact-level evaluation (Appendix C).

### v3

@@ 4.6 Role of Fact-Level Detection in Hallucination Correction
Subsequently, we evaluated the factuality of the corrected responses using the LLM-as-judge approach (Zheng et al., 2023). For each corrected sentence, we provided the external biography from Wikipedia as a reference. As elaborated in Section 4.1, this approach has been tested for annotating hallucinations and demonstrates high alignment with human annotations (Janiak et al., 2025). We instructed LLM-as-judge to return ’yes’ if the source supported the sentence, ’no’ if it was not, or ’refused’ if the LLM declined to correct the sentence (e.g., due to insufficient knowledge). We then categorized the responses into three labels: ’factual’, ’non-factual’, ’refused’.
As mentioned in Section 4.3, we utilized GPT-4o for correction and judging instead of Llama-3.1-70B-Instruct (used in detection). This choice was motivated by the challenging nature of the correction task – the model needs to correct hallucinations using only its internal knowledge, without access to external references. While the model knows hallucinated parts, it must rely on its knowledge to determine the correct information.
@@ 5 Results
This section presents the results of the experiments described in Section 4. The additional results are presented in Appendix C.
@@ 5.1 Sentence-Level Detection
Max($\mathcal{H}$) - 82.56
Mean(-logpp) - 79.20
Mean($\mathcal{H}$) - 79.02
Max(-logpp) - 78.41
@@
Tables 1 and 2 present a comparative analysis of our method against baselines. For WikiBio, FactSelfCheck-Text utilizing maxmax as an aggregation function achieves an AUC-PR score of $92.45$. It demonstrates that our approach is comparable in performance to the leading SelfCheckGPT (SCGPT) variants – Prompt ($93.60$) and NLI ($92.50$). Notably, while our method operates at a more granular level, it maintains competitive performance with a marginal decrease of $1.2\%$ compared to the best SCGPT. It is important to note that comparing our method to SelfCheckGPT at the sentence level inherently disadvantages our approach, as we operate at a more granular level of analysis. Nevertheless, our method still achieves competitive performance despite this inherent challenge.
@@
For FavaMultiSamples, FactSelfCheck-KG (Frequency-based) with maxmax aggregation achieves the highest AUC-PR score of $48.52$, outperforming all sentence-level baselines, including SelfCheckGPT (Prompt) at $46.91$. This result, together with the findings from WikiBio, highlights that the best-performing FactSelfCheck (FSC) variant depends on the dataset. On WikiBio, FSC-Text, which performs direct comparison between facts and samples, consistently achieves the highest hallucination AUC-PR ($92.45$), outperforming FSC-KG, which relies on knowledge graph comparisons. FSC-Text offers computational advantages by eliminating the knowledge graph extraction step. However, on FavaMultiSamples, the frequency-based FSC-KG variant surpasses both FSC-Text and LLM-based FSC-KG. These differences suggest that the optimal variant is influenced by dataset characteristics, such as fact density, sentence length, and text style (e.g., the prevalence of lists in FavaMultiSamples). FSC-KG may be more computationally efficient for longer samples with lower fact density due to reduced token usage, while FSC-Text is preferable for shorter or denser samples. Regarding aggregation functions, maxmax outperformed meanmean across both datasets. This makes intuitive sense – a sentence is hallucinated if any fact within it is hallucinated.
@@
An interesting side observation is that our reproduced SCGPT (Prompt) with Llama-3.1-70B-Instruct marginally surpassed the original implementation using GPT-3.5-turbo ($93.60$ vs $93.42$  [^7] The scores on GPT-3.5-turbo were obtained from the original paper (Manakul et al., 2023).).
5.2 Fact-Level Detection
Method AUC-PR
FactSelfCheck-Text 93.41
FactSelfCheck-KG (LLM-based) 92.25
FactSelfCheck-KG (Freq.-based) 87.99
SelfCheckGPT (Prompt) 86.18
RandomFact 65.79
Table 3 presents the comparative results for fact-level hallucination detection on WikiBio. FactSelfCheck-Text demonstrates superior performance with an AUC-PR score of $93.41$, followed by FactSelfCheck-KG (LLM-based) at $92.25$. The frequency-based method achieves $87.99$, while SelfCheckGPT (Prompt) scores $86.18$. These results demonstrate the effectiveness of our method in detecting hallucinations at the fact level. Furthermore, the lower performance of averaging the sentence-level scores highlights the importance of designing fact-level methods and validating our approach.
Comparing FactSelfCheck to SelfCheckGPT could be seen as unfair because the latter operates at a lower granularity than required by the evaluation task. However, this situation is analogous to the sentence-level evaluation presented in Section 5.1. In both cases, direct comparisons are not fully appropriate. The key difference is that, at the fact level, FactSelfCheck significantly outperforms SelfCheckGPT, while at the sentence level, FactSelfCheck remains competitive despite the unfavorable comparison.
@@ 6 Conclusion
We introduced FactSelfCheck, a fact-level hallucination detection approach that achieves competitive performance with existing methods while providing more interpretable insights through structured knowledge representation. By detecting hallucinations at the granular fact level rather than sentence level, our method enables more effective hallucination correction through precise identification of incorrect facts. The zero-resource nature of our approach makes it broadly applicable across diverse domains without requiring external knowledge bases or domain-specific training data. Additionally, we contributed FavaMultiSamples, a novel benchmark addressing the critical gap in evaluation datasets for sampling-based hallucination detection methods.
@@
Our study faces three primary limitations. First, we are constrained by the availability of suitable datasets. Although we contributed FavaMultiSamples, the second dataset for evaluating sampling-based methods, it is still a dataset with annotations at the sentence-level. The lack of datasets with long generated passages and annotations at the fact-level forced us to evaluate our method through aggregation rather than directly assessing our fact-level detection capabilities. Second, while the granular approach of FactSelfCheck justifies its increased complexity, the multiple LLM-based steps make it more computationally intensive compared to more straightforward methods like SelfCheckGPT. Third, our method could face challenges with very short or uninformative sentences where fact extraction may fail.
@@ Acknowledgments
This work was funded by the European Union under the Horizon Europe grant OMINO – Overcoming Multilevel INformation Overload (grant number 101086321, http://ominoproject.eu/). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. It was also co-financed with funds from the Polish Ministry of Education and Science under the programme entitled International Co-Financed Projects, grant no. 573977. This work was co-funded by the National Science Centre, Poland under CHIST-ERA Open & Re-usable Research Data & Software (grant number 2022/04/Y/ST6/00183). The infrastructure was provided by the CLARIN-PL project financed as part of the investment: "CLARIN ERIC – European Research Infrastructure Consortium: Common Language Resources and Technology Infrastructure (period: 2024-2026) funded by the Polish Ministry of Science and Higher Education (Programme: "Support for the participation of Polish scientific teams in international research infrastructure projects"), agreement number 2024/WK/01.
Appendix
This appendix provides supplementary material organized as follows:
Appendix A: FavaMultiSamples dataset creation and annotation methodology.
Appendix B: Comprehensive statistics for both evaluation datasets.
Appendix C: Additional results including complete sentence-level metrics with precision-recall curves, sample size ablation study, and evaluation of the intermediate steps of the pipeline. As well as confusion matrices and prediction statistics for the fact-level evaluation.
Appendix D: Computational complexity analysis.
Appendix E: Concrete example with knowledge graph visualization demonstrating fact-level detection advantages.
Appendix F: Enhanced prompt experiment validating comparison fairness with SelfCheckGPT.
Appendix A FavaMultiSamples Dataset
Addressing the lack of evaluation benchmarks for sampling-based hallucination detection methods, we built FavaMultiSamples upon the FAVA dataset [^8] https://huggingface.co/datasets/fava-uw/fava-data license: CC-BY-4.0 developed by Mishra et al. (2024). Prior to our contribution, WikiBio was the only available dataset for evaluating sampling-based methods, limiting evaluation of different approaches. The original FAVA dataset contains $460$ passages generated by GPT (gpt-3.5-turbo-0301) and Llama2-Chat-70B in response to diverse information-seeking prompts. Each passage was annotated by trained annotators for factual accuracy. For more details about the dataset construction and annotation process, please refer to the original paper.
@@ Appendix B Dataset Statistics
Appendix C Additional Results
C.1 Sentence-Level Detection
@@
Max($\mathcal{H}$) - 82.56 41.80 62.18
Mean(-logpp) - 79.20 44.11 61.65
Mean($\mathcal{H}$) - 79.02 49.18 64.10
Max(-logpp) - 78.41 33.59 56.00
@@
C.2 Fact-Level Detection
Pred Fact. Pred Hall.
True Fact. 951 832
True Hall. 175 3530
Table 8: Confusion matrix of FactSelfCheck-Text on fact-level evaluation.
Pred Fact. Pred Hall.
True Fact. 570 1213
True Hall. 202 3503
Table 9: Confusion matrix of SelfCheckGPT (Prompt) on fact-level evaluation.
Method Correct Incorrect
SelfCheckGPT 4073 1415
FactSelfCheck 4481 1007
Table 10: Overall prediction accuracy comparison for fact-level evaluation.
To extend the demonstration of the effectiveness of our fact-level detection approach, we present confusion matrices and prediction statistics. Table 8 shows the confusion matrix for FactSelfCheck-Text, while Table 9 presents results for SelfCheckGPT (Prompt) when adapted to fact-level prediction.
FactSelfCheck-Text correctly identifies $3530$ hallucinated facts and $951$ factual facts, achieving higher true positive rates for hallucination detection compared to SelfCheckGPT, which correctly identifies $3503$ hallucinated facts but only $570$ factual facts.
Table 10 summarizes the overall prediction accuracy of both methods. FactSelfCheck achieves $4481$ correct predictions compared to $4073$ for SelfCheckGPT, representing an increase of $408$ correct predictions ($10.02\%$ improvement).
C.3 Effect of Sample Size on Detection Performance
@@
Figure 5 shows a different pattern for FavaMultiSamples, where hallucination detection performance decreases with more samples. This occurs because methods with few samples tend to overestimate hallucination scores, interpreting normal variations as potential hallucinations. As sample size increases, methods become more conservative in their scoring, leading to better calibration. This pattern likely stems from FavaMultiSamples having shorter sentences with fewer facts compared to WikiBio (see Section C.4).
@@
C.4 Evaluation of Intermediate Steps
# entities/ # relations/ # facts/
@@ C.4 Evaluation of Intermediate Steps
Appendix D Computational complexity
@@
Appendix E Case Study of FactSelfCheck vs SelfCheckGPT
@@
Table 13: Comparison of fact-level FactSelfCheck with sentence-level SelfCheckGPT. An external Wikipedia biography is provided to analyse the correctness of the methods. The red value indicates hallucinations, and the green value indicates factual correctness. The facts were classified using a threshold of $0.3$ utilizing FactSelfCheck, and the sentences were classified using a threshold of $0.75$ with SelfCheckGPT. These thresholds achieved the highest F1-scores in fact-level (Appendix C.2) and sentence-level (Section 5.1) evaluation, respectively.
Figure 6: Example of a knowledge graph extracted from a response. Edge width represents the hallucination score for each fact, with red edges indicating hallucinated facts and green edges indicating correct facts. Facts were classified using a threshold of $0.3$, which achieved the highest F1-score in the fact-level evaluation (C.2).
Appendix F SelfCheckGPT with Enhanced Prompt